# Parallel Token Swapping for Qubit Routing

### v1

@@ 4 Cycle graphs
Let $e=(i,i+1)$ be an edge in $E$. Let $P_{e}$ be the line graph obtained by deleting from $G$ the edge diametrically opposite to edge $E$. We say that an edge $e=(i,i+1)\in E$ is reasonable with respect to a configuration $C$ if the tokens $C(i)$ and $C(i+1)$ on the line graph $P_{e}$ occur in order opposite to what their order on this path would be in the identity configuration. As an example, consider the edge $e=(r,r+1)$. The diametrically opposite edge is $(2r,1)$ and deleting this we obtain the path $(1,2,3,\ldots,2r)$. Thus, edge $e=(r,r+1)$ is reasonable with respect to a configuration $C$ if and only if $C(r)>C(r+1)$. Consider a modification of Algorithm 1 that alternates between swapping all reasonable odd edges and all reasonable even edges. Call this Algorithm 2, specified below.
@@
Let $OPT=(M_{1},M_{2},\ldots,M_{k^{*}})$ be an optimal solution. By Observation 9, we know that $k_{2}\leq 2r$. Hence if $k^{*}\geq r$, then Algorithm 3 outputs a solution of size at most $2OPT$. So assume that $k^{*}<n$.
@@
Let $OPT=(M_{1},M_{2},\ldots,M_{k^{*}})$ be a reasonable solution. We can set $M^{\prime}_{1}=M_{1}\cap E_{1}$, $M^{\prime}_{2}=(M_{1}\cup M_{2})\cap E_{2}$, $M^{\prime}_{3}=(M_{2}\cup M_{3})\cap E_{1}$, $\ldots$, $M^{\prime}_{k^{*}+1}=M_{k^{*}}\cap E_{1}$. $\Box$
@@
Let $G=(V,E)$ be an odd-cycle graph with vertices $V=(1,2,\ldots,n=2r+1)$ and edges $E=\{(i,i+1):1\leq i\leq n\}$ with the understanding that $n+1=1$. The definition of a reasonable swap from our discussions on even-cycles needs to be modified as the diametrically opposite edge is not well-defined (there are two such edges). We fix diametrically opposite to mean the first of the two in the clockwise direction. Additionally, we define $E_{1}={(1,2),(3,4),\ldots,(2r-1,2r)}$, $E_{3}=(2r+1,1)$ and $E_{2}=E\setminus(E_{1}\cup E_{3})$. We define a $(E_{1},E_{3},E_{2},E_{3})$-alternating solution to be one where $M_{4k+1}\subseteq E_{1},M_{4k+3}\subseteq E_{2}$ and $M_{2k}\subseteq E_{3}$.
If the optimal solution $M_{1},\ldots,M_{k^{*}}$ has a token which walks half-way around the cycle, then Algorithm 1 on the input $G^{\prime}=(V,E\setminus\{(2r+1,1)\})$ outputs a solution of size at most $12k^{*}+$ 1. If not, we can covert $OPT$ to an $(E_{1},E_{3},E_{2},E_{3})$-alternating solution such that no token moves half-way across the cycle and of size at most $12k^{*}+$ 1 as follows. Set $M^{\prime}_{1}=M_{1}\cap E_{1}$, $M^{\prime}_{2}=M_{1}\cap E_{3}$, $M^{\prime}_{3}=(M_{1}\cup M_{2})\cap E_{2}$, $M^{\prime}_{4}=M_{2}\cap E_{3}$, $M^{\prime}_{5}=(M_{2}\cup M_{3})\cap E_{1}$ and so on. Next, we can modify Algorithm 2 so that in iteration $14k+$ 1, we make all reasonable swaps from the edge set $E_{1}$, in iteration $34k+$ 3, we make all reasonable swaps from the edge set $E_{2}$, and in iteration $2k$ we make all reasonable swaps from the edge set $E_{3}$. As in the case of even-cycles, we argue that this algorithm outputs the best $(E_{1},E_{3},E_{2},E_{3})$-alternating solution such that no token moves half-way across the cycle. This completes the proof. ∎
@@ 5 Subdivided star graphs
In this section, we consider the case when the input graph is a subdivided star. A subdivided star is a graph consisting of any number of paths (that we call branches) all joined at a single endpoint. Suppose we have a subdivided star graph with branches $1,\dots,h$, and each branch $b$ contains $n_{b}$ vertices and edges. We label the center vertex $v_{0}$, the vertices in branch $b$ as $v_{b,1},\dots,v_{b,n_{b}}$, and the edges of branch $b$ as $(v_{b,1},v_{b,2}),\dots,(v_{b,n_{b}-1}),(v_{b,n})$ and $(v_{0},v_{b,1})$. To make our notation more convenient, for every branch $b$, we let $v_{b,0}=v_{0}$ be an equivalent name for the center vertex, and we let the set $B_{b}$ be both the set of vertices in the branch as well as the tokens that belong in the branch.
@@
Our algorithm proceeds in three phases. In the Phase 1, we sort every branch $b$ so that any token that does not belong on $B_{b}$ is closer to $v_{0}$ than any token that does belong on $B_{b}$. (Let us call a configuration branch-sorted if it satisfies this condition.) In Phase 2, we route all tokens to the correct branch in a greedy manner. While doing this, we make sure that the center token $v_{0}$ returns to its destination as infrequently as possible. Finally, in Phase 3, we use Algorithm 1 to sort the branches independently to reach the identity configuration.
@@
In any solution, at most one token gets moved onto its correct branch in each matching, so the size of any optimal solution is at least the number of tokens on the wrong branch. In Phase 2, every iteration moves one token onto its correct branch unless $v_{0}$ is at the center. Token $v_{0}$ can only arrive at the center once per branch that contains a token which needs to switch branches. This number is bounded by $\min\{OPT,h\}$, and so the total number of iterations is at most $\min\{OPT,h\}$ more than the number of tokens that must switch branches. Therefore, we can bound Phase 2 by $OPT+h$. $\Box$
@@
Define $E_{r}=\{(v_{b,i-1},v_{b,i}:1\leq i\leq r\}$ be edges within a radius $r$ from the center vertex $v_{0}$. Let $M_{i,1}=M_{i}^{*}\cap E_{k_{1}-i+1}$ and $M_{i,2}=M_{i}^{*}\setminus M_{i,1}$. Then we claim that the sequence of matchings $\mathcal{M}=M_{1,1},\dots,M_{k_{1},1},M_{1,2},\dots,M_{k_{1},2}$ applied to $C_{0}$ arrives at the same configuration as $M_{k_{1}}^{*}\dots M_{1}^{*}C_{0}$. Note that since the radius of our edge set decreases by one for each successive matching, swaps made outside the radius (those in $M_{1,2},\dots,M_{k_{1},2}$) cannot affect the tokens being swapped in $M_{1,1},\dots,M_{k_{1},1}$. Therefore, we can defer swaps $M_{i,2}$ until after all swaps $M_{i,1}$.
@@
Let $\hat{C}$ be the configuration that Algorithm 4 arrives at after Phase 2, and let $C^{\prime}$ be the configuration $M_{k_{1},1}\dots M_{1,1}C_{0}$. Then $\hat{C}$ can be routed to $C^{\prime}$ in at most $k_{1}$ steps: for branch $b$, either $k_{1}\geq n_{b}$, or tokens further than $k_{1}$ from $v_{0}$ are all on the correct branch. In the latter case, nothing in Phase 1 or Phase 2 of Algorithm 4 nor in $M_{1,1},\dots,M_{k_{1},1}$ involved any of these tokens, so $\hat{C}$ and $C^{\prime}$ only differ in a radius $k_{1}$ from $v_{0}$. Therefore, Algorithm 1 can route one to the other in at most $k_{1}$ steps. From $C^{\prime}$, we can arrive at the identity configuration in $k_{1}+k_{2}$ more matchings (using $M_{1,2},\dots,M_{k_{1},2},M_{k_{1}+1},\dots,M_{k_{1}+k_{2}}$). Thus, there is a solution that can route from $\hat{C}$ to $Id$ in at most $2k_{1}+k_{2}\leq 2OPT$ matchings. Since we use Algorithm 1, we are guaranteed to take no more than $12OPT+$ 1 iterations in Phase 3. $\Box$
@@ 6 Grid graphs
The first algorithm simply considers a spanning subgraph $G^{\prime}$ of $G$ such that $G^{\prime}$ is a line graph and uses Algorithm 1 on $G^{\prime}$ to route the tokens to their correct positions. We show that the algorithm outputs a solution of size at most $12OPT+$ 1 when $h=2$. Such grid graphs are also called ladder graphs.
@@
1:$E_{1}\leftarrow\{\bigl{(}(a,b),(a+1,b)\bigr{)}:1\leq a\leq h-1,1\leq b\leq n\}\cup\{\bigl{(}(h,b),(h,b+1)\bigr{)}:b=1,3,5,\ldots\}\cup\{\bigl{(}(1,b),(1,b+1)\bigr{)}:b=2,4,6,\ldots\}$.
2:$M_{1},\ldots,M_{k}\leftarrow$ Output of Algorithm 1 on graph $G^{\prime}=(V,E_{1})$ and configuration $C$.
@@
Algorithm 5 outputs a solution of size at most $12OPT+$ 1 for (PTS) on $2\times n$ grid graphs.
@@
Let $M^{*}_{1},\ldots,M^{*}_{k^{*}}$ be an optimal solution. The graph $G^{\prime}=(V,E_{1})$ in the algorithm is formed by picking every horizontal edge and alternating between the vertical edges in the first and last columns. This creates a line graph. Let $A=\{\bigl{(}(1,b),(2,b)\bigr{)}:1\leq b\leq n\}$ be all the horizontal edges and $B=\{\bigl{(}(2,b),(2,b+1)\bigr{)}:b=1,3,5,\ldots\}\cup\{\bigl{(}(1,b),(1,b+1)\bigr{)}:b=2,4,6,\ldots\}$ be all the vertical edges in $E_{1}$. Let $E\setminus E_{1}=E_{2}$. We say that a solution $M_{1},\ldots,M_{k}$ is $(A,B)$-alternating if $M_{2i+1}\subseteq A$ and $M_{2i}\subseteq B$.
We show that there exists an $(A,B)$-alternating solution that with size at most $12k^{*}+$ 1. This implies that Algorithm 5 outputs a solution of size at most $k^{*}+1$ since Algorithm 1 outputs the best $(A,B)$-alternating solution on line graphs (This is precisely Theorem 7 in [12]).
We begin by expanding the optimal solution into a new feasible solution as follows: $M^{*}_{1}\cap E_{1},M^{*}_{1}\cap E_{2},M^{*}_{2}\cap E_{2},M^{*}_{2}\cap E_{1},M^{*}_{3}\cap E_{1},\ldots$. Observe that the set of edges $E_{2}$ is a matching in itself and so any consecutive matchings $M^{*}_{i}\cap E_{2}$ and $M^{*}_{i+1}\cap E_{2}$ can be combined into a single matching. Hence, we obtain the following feasible solution, $M^{*}_{1}\cap E_{1},(M^{*}_{1}\cup M^{*}_{2})\cap E_{2},M^{*}_{2}\cap E_{1},M^{*}_{3}\cap(M^{*}_{3}\cup M^{*}_{4})\cap E_{2},\ldots$, of size at most $k^{*}+1$. We now show how to simulate a matching in $E_{2}$ using only the edges in $E_{1}$.
@@
Let $M\subseteq E_{2}$ and $C$ be any configuration. Then, there exists three matchings $M_{1},M_{2},M_{3}$ such that $M_{3}M_{2}M_{1}C=MC$, and $M_{1}\subseteq A,M_{2}\subseteq B$ and $M_{3}\subseteq A$.
@@
Consider the case when $M=E_{2}$. Then setting $M_{1}=A,M_{2}=B$ and $M_{3}=A$ proves the result. Now if $M\neq E_{2}$, we can reduce to the previous case by splitting the ladder graph into disjoint smaller ladder graphs, where the above case applies. These can be routed in parallel. For instance, if $M=\{\bigl{(}(1,1),(1,2)\bigr{)},\bigl{(}(2,4),(2,5)\bigr{)}\}$. Then, the smaller ladder graphs are obtained by taking the vertices in the first two rows, and by taking the vertices in rows 4 and 5. Then, $M_{1}=M_{3}=\{\bigl{(}(1,1),(2,1)\bigr{)},\bigl{(}(2,1),(2,2)\bigr{)},\bigl{(}(1,4),(2,4)\bigr{)},\bigl{(}(1,5),(2,5)\bigr{)}\}$ and $M_{2}=\{\bigl{(}(2,1),(2,2)\bigr{)},\bigl{(}(1,4),(1,5)\bigr{)}\}$. $\Box$
We apply Claim G to the feasible solution of size $k^{*}+1$ from earlier to replace matchings in $E_{2}$ with three matchings in $E_{1}$. Furthermore, we split a matching $M_{i}\cap E_{1}$ into either $M_{i}\cap A,M_{i}\cap B$ or $M_{i}\cap B,M_{i}\cap A$. We use $A_{i}$ to denote a matching that lies in the edge set $A$ and $B_{i}$ to denote a matching that lies in the edge set $B$. Then, we can change the feasible solution $M^{*}_{1}\cap E_{1},(M^{*}_{1}\cup M^{*}_{2})\cap E_{2},M^{*}_{2}\cap E_{1},M^{*}_{3}\cap(M^{*}_{3}\cup M^{*}_{4})\cap E_{2},\ldots$ from earlier into a feasible solution of the following form, $(B_{1},A_{1}),(A_{2},B_{2},A_{3}),(A_{4},B_{4}),(B_{5},A_{5}),(A_{6},B_{6},A_{7}),\ldots$. But now again, since edge sets $A$ and $B$ are matchings by themselves, we can combine consecutive matchings $A_{i}$ and $A_{i+1}$ into a single matching and consecutive matchings $B_{i}$ and $B_{i+1}$ into a single matching. Finally, we obtain the following feasible solution, $B_{1}^{\prime},A_{1}^{\prime},B_{2}^{\prime},A_{2}^{\prime},B_{3}^{\prime},A_{3}^{\prime},B_{4}^{\prime},A_{4}^{\prime},\ldots$ of size at most $12k^{*}+$ 1. Thus, we have exhibited an $(A,B)$-alternating solution of size at most $12k^{*}+$ 1, completing the proof of the theorem. ∎
@@
Algorithm 6 works in three phases. We bound the number of matchings in each of the phases separately. In the first phase, the algorithm routes the configuration $C$ into the configuration $C_{1}$ using Algorithm 1 on the rows. Due to Observation 9, this takes at most $h$ matchings. The second phase routes the configuration $C_{1}$ into the configuration $C_{2}$ using Algorithm 1 on the columns. Let $d_{max}$ be the maximum geodesic distance between $C^{-1}(t)$ and $Id^{-1}(t)$ in the graph $G$ for any token $t$. Let $d^{\prime}$ be the maximum geodesic distance between $C_{1}^{-1}(t)$ and $C_{2}^{-1}(t)$ in the graph $G$ for any token $t$. The geodesic distance between $C_{1}^{-1}(t)$ and $C_{2}^{-1}(t)$ for any token $t$ corresponds to the vertical distance between $C^{-1}(t)$ and $Id^{-1}(t)$. Hence, $d^{\prime}\leq d_{max}$.

### v2

@@ 4 Cycle graphs
Let $e=(i,i+1)$ be an edge in $E$. Define pairs of diametrically opposite edges to be $\left\{(i,i+1),(r+i,r+i+1)\right\}$ for $i=1,2,\ldots,r-1$, and $\left\{(r,r+1),(2r,1)\right\}$ (see Figure 1). Let $P_{e}$ be the line graph obtained by deleting from $G$ the edge diametrically opposite to edge $e$. We say that an edge $e=(i,i+1)\in E$ is reasonable with respect to a configuration $C$ if the tokens $C(i)$ and $C(i+1)$ on the line graph $P_{e}$ occur in order opposite to what their order on this path would be in the identity configuration. As an example, consider the edge $e=(r,r+1)$. The diametrically opposite edge is $(2r,1)$ and deleting this we obtain the path $(1,2,3,\ldots,2r)$. Thus, edge $e=(r,r+1)$ is reasonable with respect to a configuration $C$ if and only if $C(r)>C(r+1)$. Consider a modification of Algorithm 1 that alternates between swapping all reasonable odd edges and all reasonable even edges. Call this Algorithm 2, specified below.
@@
Let $OPT=(M_{1},M_{2},\ldots,M_{k^{*}})$ be an optimal solution. By Observation 10, we know that $k_{2}\leq 2r$. Hence if $k^{*}\geq r$, then Algorithm 3 outputs a solution of size at most $2OPT$. So assume that $k^{*}<r$.
@@
Let $OPT=(M_{1},M_{2},\ldots,M_{k^{*}})$ be a reasonable solution. We can set $M^{\prime}_{1}=M_{1}\cap E_{1}$, $M^{\prime}_{2}=(M_{1}\cup M_{2})\cap E_{2}$, $M^{\prime}_{3}=(M_{2}\cup M_{3})\cap E_{1}$, $\ldots$, $M^{\prime}_{k^{*}+1}$, where $M^{\prime}_{k^{*}+1}=M_{k^{*}}\cap E_{1}$ if $k^{*}$ is even and $M^{\prime}_{k^{*}+1}=M_{k^{*}}\cap E_{2}$ if $k^{*}$ is odd. Recall from Remark 7 that $M_{i}\cap M_{i+1}=\emptyset$. Hence, the new sequence $(M^{\prime}_{1},\ldots,M^{\prime}_{k^{*}+1})$ is exactly mimicking the original sequence $M_{1},\ldots,M_{k^{*}}$ implying that the new sequence is an $(E_{1},E_{2})$-alternating reasonable solution of size $k^{*}+1$ such that no token is routed half-way around the cycle. $\Box$
@@
Let $G=(V,E)$ be an odd-cycle graph with vertices $V=(1,2,\ldots,n=2r+1)$ and edges $E=\{(i,i+1):1\leq i\leq n\}$ with the understanding that $n+1=1$. Define pairs of diametrically opposite edges to be $\left\{(i,i+1),(r+i,r+i+1)\right\}$ for $i=1,2,\ldots,r$, and $\left\{(r+1,r+2),(2r,1)\right\}$. We can then define reasonable swaps exactly as in the case of even cycles. Additionally, we define $E_{1}={(1,2),(3,4),\ldots,(2r-1,2r)}$, $E_{3}=(2r+1,1)$ and $E_{2}=E\setminus(E_{1}\cup E_{3})$. We define a $(E_{1},E_{3},E_{2},E_{3})$-alternating solution to be one where $M_{4k+1}\subseteq E_{1},M_{4k+3}\subseteq E_{2}$ and $M_{2k}\subseteq E_{3}$.
If the optimal solution $M_{1},\ldots,M_{k^{*}}$ has a token which walks half-way around the cycle, then Algorithm 1 on the input $G^{\prime}=(V,E\setminus\{(2r+1,1)\})$ outputs a solution of size at most $12k^{*}+$ 1. If not, we can covert $OPT$ to an $(E_{1},E_{3},E_{2},E_{3})$-alternating solution such that no token moves half-way across the cycle and of size at most $12k^{*}+$ 1 as follows. Set $M^{\prime}_{1}=M_{1}\cap E_{1}$, $M^{\prime}_{2}=M_{1}\cap E_{3}$, $M^{\prime}_{3}=(M_{1}\cup M_{2})\cap E_{2}$, $M^{\prime}_{4}=M_{2}\cap E_{3}$, $M^{\prime}_{5}=(M_{2}\cup M_{3})\cap E_{1}$ and so on up to $M^{\prime}_{2k^{*}}=M_{k^{*}}\cap E_{3}$ and $M^{\prime}_{2k^{*}+1}=M_{k^{*}}\cap E_{1}$ if $k^{*}$ is even and $M^{\prime}_{2k^{*}+1}=M_{k^{*}}\cap E_{2}$ if $k^{*}$ is odd. Next, we can modify Algorithm 2 so that in iteration $14k+$ 1, we make all reasonable swaps from the edge set $E_{1}$, in iteration $34k+$ 3, we make all reasonable swaps from the edge set $E_{2}$, and in iteration $2k$ we make all reasonable swaps from the edge set $E_{3}$. As in the case of even-cycles, we argue that this algorithm outputs the best $(E_{1},E_{3},E_{2},E_{3})$-alternating solution such that no token moves half-way across the cycle. This completes the proof. ∎
@@ 5 Subdivided star graphs
In this section, we consider the case when the input graph is a subdivided star. A subdivided star is a graph consisting of any number of paths (that we call branches) all joined at a single endpoint. Suppose we have a subdivided star graph with branches $1,\dots,h$, and each branch $b$ contains $n_{b}$ vertices and edges. We label the center vertex $v_{0}$, the vertices in branch $b$ as $v_{b,1},\dots,v_{b,n_{b}}$, and the edges of branch $b$ as $(v_{b,1},v_{b,2}),\dots,(v_{b,n_{b}-1},v_{b,n})$ and $(v_{0},v_{b,1})$. To make our notation more convenient, for every branch $b$, we let $v_{b,0}=v_{0}$ be an equivalent name for the center vertex, and we let the set $B_{b}$ be both the set of vertices in the branch as well as the tokens that belong in the branch.
@@
Our algorithm proceeds in three phases. In Phase 1, we sort every branch $b$ so that any token that does not belong on $B_{b}$ is closer to $v_{0}$ than any token that does belong on $B_{b}$. (Let us call a configuration branch-sorted if it satisfies this condition.) In Phase 2, we route all tokens to the correct branch in a greedy manner. While doing this, we make sure that the center token $v_{0}$ returns to its destination as infrequently as possible. Finally, in Phase 3, we use Algorithm 1 to sort the branches independently to reach the identity configuration.
@@
In any solution, at most one token gets moved onto its correct branch in each matching, so the size of any optimal solution is at least the number of tokens on the wrong branch. In Phase 2, every iteration moves one token onto its correct branch unless $v_{0}$ is at the center. Token $v_{0}$ can only arrive at the center once per branch that contains a token which needs to switch branches. This number is bounded by $\min\{OPT,h\}$, and so the total number of iterations is at most $\min\{OPT,h\}$ more than the number of tokens that must switch branches. Therefore, we can bound Phase 2 by $OPT+\min\{OPT,h\}$. $\Box$
@@
Define $E_{r}=\{(v_{b,i-1},v_{b,i}):1\leq i\leq r,i\leq n_{b}\}$ to be edges within a radius $r$ from the center vertex $v_{0}$. Let $M_{i,1}=M_{i}^{*}\cap E_{k_{1}-i+1}$ and $M_{i,2}=M_{i}^{*}\setminus M_{i,1}$. Then we claim that the sequence of matchings $\mathcal{M}=M_{1,1},\dots,M_{k_{1},1},M_{1,2},\dots,M_{k_{1},2}$ applied to $C_{0}$ arrives at the same configuration as $M_{k_{1}}^{*}\dots M_{1}^{*}C_{0}$. Note that since the radius of our edge set decreases by one for each successive matching, swaps made outside the radius (those in $M_{1,2},\dots,M_{k_{1},2}$) cannot affect the tokens being swapped in $M_{1,1},\dots,M_{k_{1},1}$. Therefore, we can defer swaps $M_{i,2}$ until after all swaps $M_{i,1}$.
@@
Let $\hat{C}$ be the configuration that Algorithm 4 arrives at after Phase 2, and let $C^{\prime}$ be the configuration $M_{k_{1},1}\dots M_{1,1}C_{0}$. Then $\hat{C}$ can be routed to $C^{\prime}$ in at most $k_{1}$ steps: for branch $b$, either $k_{1}\geq n_{b}$, or tokens further than $k_{1}$ from $v_{0}$ are all on the correct branch. In the latter case, nothing in Phase 1 or Phase 2 of Algorithm 4 nor in $M_{1,1},\dots,M_{k_{1},1}$ involved any of these tokens, so $\hat{C}$ and $C^{\prime}$ only differ in a radius $k_{1}$ from $v_{0}$. Therefore, Algorithm 1 can route one to the other in at most $k_{1}$ steps. From $C^{\prime}$, we can arrive at the identity configuration in $k_{1}+k_{2}$ more matchings (using $M_{1,2},\dots,M_{k_{1},2},M^{*}_{k_{1}+1},\dots,M^{*}_{k_{1}+k_{2}}$). Thus, there is a solution that can route from $\hat{C}$ to $Id$ in at most $2k_{1}+k_{2}\leq 2OPT$ matchings. Since we use Algorithm 1, we are guaranteed to take no more than $12OPT+$ 1 iterations in Phase 3. $\Box$
@@ 6 Grid graphs
The first algorithm simply considers a spanning subgraph $G^{\prime}$ of $G$ such that $G^{\prime}$ is a line graph and uses Algorithm 1 on $G^{\prime}$ to route the tokens to their correct positions. We show that the algorithm outputs a solution of size at most $22OPT+$ 2 when $h=2$. Such grid graphs are also called ladder graphs.
[…]
Figure 3: The partitioning of the edge set $E$ into $F_{2},A$ and $B$ in ladder graphs. $F_{1}=A\cup B$.
@@
1:$F_{1}\leftarrow\{\bigl{(}(a,b),(a+1,b)\bigr{)}:1\leq a\leq h-1,1\leq b\leq n\}\cup\{\bigl{(}(h,b),(h,b+1)\bigr{)}:b=1,3,5,\ldots\}\cup\{\bigl{(}(1,b),(1,b+1)\bigr{)}:b=2,4,6,\ldots\}$.
2:$M_{1},\ldots,M_{k}\leftarrow$ Output of Algorithm 1 on graph $G^{\prime}=(V,F_{1})$ and configuration $C$.
@@
Algorithm 5 outputs a solution of size at most $22OPT+$ 2 for (PTS) on $2\times n$ grid graphs.
@@
Let $M^{*}_{1},\ldots,M^{*}_{k^{*}}$ be an optimal solution. The graph $G^{\prime}=(V,F_{1})$ in the algorithm is formed by picking every horizontal edge and alternating between the vertical edges in the first and last columns. This creates a line graph. Let $A=\{\bigl{(}(1,b),(2,b)\bigr{)}:1\leq b\leq n\}$ be all the horizontal edges and $B=\{\bigl{(}(2,b),(2,b+1)\bigr{)}:b=1,3,5,\ldots\}\cup\{\bigl{(}(1,b),(1,b+1)\bigr{)}:b=2,4,6,\ldots\}$ be all the vertical edges in $F_{1}$. Let $E\setminus F_{1}=F_{2}$. We say that a solution $M_{1},\ldots,M_{k}$ is $(A,B)$-alternating if $M_{2i+1}\subseteq A$ and $M_{2i}\subseteq B$.
We show that there exists an $(A,B)$-alternating solution with size at most $22k^{*}+$ 2. This implies that Algorithm 5 outputs a solution of size at most $22k^{*}+$ 2 since Algorithm 1 outputs the best $(A,B)$-alternating solution on line graphs (This is precisely Theorem 7 in [14]).
We begin by expanding the optimal solution into a new feasible solution as follows: $M^{*}_{1}\cap F_{1},M^{*}_{1}\cap F_{2},M^{*}_{2}\cap F_{2},M^{*}_{2}\cap F_{1},M^{*}_{3}\cap F_{1},\ldots$. Observe that the set of edges $F_{2}$ is a matching in itself and so any consecutive matchings $M^{*}_{i}\cap F_{2}$ and $M^{*}_{i+1}\cap F_{2}$ can be combined into a single matching. Hence, we obtain the following feasible solution, $M^{*}_{1}\cap F_{1},(M^{*}_{1}\cup M^{*}_{2})\cap F_{2},M^{*}_{2}\cap F_{1},M^{*}_{3}\cap F_{1},(M^{*}_{3}\cup M^{*}_{4})\cap F_{2},\ldots$ up to $(M^{*}_{k^{*}-1}\cup M^{*}_{k^{*}})\cap F_{2},M^{*}_{k^{*}}\cap F_{1}$ if $k^{*}$ is even and $M^{*}_{k^{*}}\cap F_{1},M^{*}_{k^{*}}\cap F_{2}$ if $k^{*}$ is odd. In the above transformation, we are taking two consecutive matchings $M^{*}_{2i-1},M^{*}_{2i}$ and replacing them by three matchings $M^{*}_{2i-1}\cap F_{1},(M^{*}_{2i-1}\cup M^{*}_{2i})\cap F_{2},M^{*}_{2i}\cap F_{1}$. We now show how to simulate a matching in $F_{2}$ using only the edges in $F_{1}$.
@@
Let $M\subseteq F_{2}$ and $C$ be any configuration. Then, there exist three matchings $M_{1},M_{2},M_{3}$ such that $M_{3}M_{2}M_{1}C=MC$, and $M_{1}\subseteq A,M_{2}\subseteq B$ and $M_{3}\subseteq A$.
@@
Consider the case when $M=F_{2}$. Then setting $M_{1}=A,M_{2}=B$ and $M_{3}=A$ proves the result (see Figure 4). Now if $M\neq F_{2}$, we can reduce to the previous case by splitting the ladder graph into disjoint smaller ladder graphs, where the above case applies. These can be routed in parallel. For instance, if $M=\{\bigl{(}(1,1),(1,2)\bigr{)},\bigl{(}(2,4),(2,5)\bigr{)}\}$. Then, the smaller ladder graphs are obtained by taking the vertices in the first two rows, and by taking the vertices in rows 4 and 5. Then, $M_{1}=M_{3}=\{\bigl{(}(1,1),(2,1)\bigr{)},\bigl{(}(2,1),(2,2)\bigr{)},\bigl{(}(1,4),(2,4)\bigr{)},\bigl{(}(1,5),(2,5)\bigr{)}\}$ and $M_{2}=\{\bigl{(}(2,1),(2,2)\bigr{)},\bigl{(}(1,4),(1,5)\bigr{)}\}$. $\Box$
We apply Claim G to the feasible solution from earlier to replace matchings in $F_{2}$ with three matchings in $F_{1}$. Furthermore, we split a matching $M_{i}\cap F_{1}$ into either $M_{i}\cap A,M_{i}\cap B$ or $M_{i}\cap B,M_{i}\cap A$. We use $A_{i}$ to denote a matching that lies in the edge set $A$ and $B_{i}$ to denote a matching that lies in the edge set $B$. Then, we can change the feasible solution $M^{*}_{1}\cap F_{1},(M^{*}_{1}\cup M^{*}_{2})\cap F_{2},M^{*}_{2}\cap F_{1},M^{*}_{3}\cap F_{1},(M^{*}_{3}\cup M^{*}_{4})\cap F_{2},\ldots$ from earlier into a feasible solution of the following form, $(B_{1},A_{1}),(A_{2},B_{2},A_{3}),(A_{4},B_{3}),(B_{4},A_{5}),(A_{6},B_{5},A_{7}),\ldots$. In the above transformation, we are essentially taking two consecutive original matchings $M^{*}_{2i-1},M^{*}_{2i}$ and replacing them with $(B_{3i-2},A_{4i-3}),(A_{4i-2},B_{3i-1},A_{4i-1}),(A_{4i},B_{3i})$. But now again, since edge sets $A$ and $B$ are matchings by themselves, we can combine consecutive matchings $A_{i}$ and $A_{i+1}$ into a single matching. Hence, we can replace two consecutive original matchings $M^{*}_{2i-1},M^{*}_{2i}$ with the five matchings $(B_{3i-2},A^{\prime}_{2i-1},B_{3i-1},A^{\prime}_{2i},B_{3i})$ where $A^{\prime}_{2i-1}=A_{4i-3}\cup A_{4i-2}$ and $A^{\prime}_{2i}=A_{4i-1}\cup A_{4i}$. Note that the ends of these sequences are both matchings in $B$ and so we can combine $B_{3i-2}$ and $B_{3i-3}$ into a single matching and $B_{3i},B_{3i+1}$ into a single matching. This will give us a sequence $B_{1}^{\prime},A_{1}^{\prime},B_{2}^{\prime},A_{2}^{\prime},\cdots$ where $B^{\prime}_{2i-1}=B_{3i-3}\cup B_{3i-2}$ and $B^{\prime}_{2i}=B_{3i-1}$ and $B^{\prime}_{2i+1}=B_{3i}\cup B_{3i+1}$. If $k^{*}$ is even, the tail of this sequence ends with $A_{k^{*}},B_{k^{*}+1}$ and has a length of $12k^{*}+$ 1. On the other hand, if $k^{*}$ is odd, then the tail of this sequence will end in $B_{k^{*}+1},A_{k^{*}+1}$ and has a length of $22k^{*}+$ 2. Thus, we have exhibited an $(A,B)$-alternating solution of size at most $22k^{*}+$ 2, completing the proof of the theorem. ∎
[…]
$F_{2}$
$v_{1}$$t_{1}$$v_{4}$$t_{4}$$v_{5}$$t_{5}$$v_{8}$$t_{8}$$v_{2}$$t_{2}$$v_{3}$$t_{3}$$v_{6}$$t_{6}$$v_{7}$$t_{7}$
(a) Routing using the matching $F_{2}$
$v_{1}$$t_{4}$$v_{4}$$t_{1}$$v_{5}$$t_{8}$$v_{8}$$t_{5}$$v_{2}$$t_{2}$$v_{3}$$t_{6}$$v_{6}$$t_{3}$$v_{7}$$t_{7}$
$A$
$v_{1}$$t_{2}$$v_{4}$$t_{6}$$v_{5}$$t_{3}$$v_{8}$$t_{7}$$v_{2}$$t_{4}$$v_{3}$$t_{1}$$v_{6}$$t_{8}$$v_{7}$$t_{5}$
$B$
$v_{1}$$t_{2}$$v_{4}$$t_{3}$$v_{5}$$t_{6}$$v_{8}$$t_{7}$$v_{2}$$t_{1}$$v_{3}$$t_{4}$$v_{6}$$t_{5}$$v_{7}$$t_{8}$
$A$
$v_{1}$$t_{1}$$v_{4}$$t_{4}$$v_{5}$$t_{5}$$v_{8}$$t_{8}$$v_{2}$$t_{2}$$v_{3}$$t_{3}$$v_{6}$$t_{6}$$v_{7}$$t_{7}$
(b) Routing using the three matchings $A,B$ and $A$
Figure 4: Simulating the routing from $F_{2}$ using the three matchings $A,B$ and $A$.
@@
Algorithm 6 works in three phases. We bound the number of matchings in each of the phases separately. In the first phase, the algorithm routes the configuration $C$ into the configuration $C_{1}$ using Algorithm 1 on the rows. Due to Observation 10, this takes at most $h$ matchings. The second phase routes the configuration $C_{1}$ into the configuration $C_{2}$ using Algorithm 1 on the columns. Let $d_{max}$ be the maximum geodesic distance between $C^{-1}(t)$ and $Id^{-1}(t)$ in the graph $G$ for any token $t$. Let $d^{\prime}$ be the maximum geodesic distance between $C_{1}^{-1}(t)$ and $C_{2}^{-1}(t)$ in the graph $G$ for any token $t$. The geodesic distance between $C_{1}^{-1}(t)$ and $C_{2}^{-1}(t)$ for any token $t$ corresponds to the vertical distance between $C^{-1}(t)$ and $Id^{-1}(t)$. Hence, $d^{\prime}\leq d_{max}$.